OpenStudy (angelab97):

If a+b= -8 and a^2+b^2=50 what is ab? What's a shortcut to solving this problem?

OpenStudy (aum):

\(\large (a+b)^2 = a^2 + 2ab + b^2 = (a^2+b^2) + 2ab\) Substitute the values: (a+b) = -8; a^2 + b^2 = 50 \(\large (-8)^2 = (50) + 2ab\). Solve for ab.

OpenStudy (angelab97):

thank you so much :)

OpenStudy (aum):

you are welcome.